# Satellite image classification with neural quantum kernels

### v1

@@ Abstract
A practical application of quantum machine learning in real-world scenarios in the short term remains elusive, despite significant theoretical efforts. Image classification, a common task for classical models, has been used to benchmark quantum algorithms with simple datasets, but only few studies have tackled complex real-data classification challenges. In this work, we address such a gap by focusing on the classification of satellite images, a task of particular interest to the earth observation (EO) industry. We first preprocess the selected intrincate dataset by reducing its dimensionality. Subsequently, we employ neural quantum kernels (NQKs) -embedding quantum kernels (EQKs) constructed from trained quantum neural networks (QNNs)- to classify images which include solar panels. We explore both $1$-to-$n$ and $n$-to-$n$ NQKs. In the former, parameters from a single-qubit QNN’s training construct an $n$-qubit EQK achieving a mean test accuracy over 86% with three features. In the latter, we iteratively train an $n$-qubit QNN to ensure scalability, using the resultant architecture to directly form an $n$-qubit EQK. In this case, a test accuracy over 88% is obtained for three features and 8 qubits. Additionally, we show that the results are robust against a suboptimal training of the QNN.
@@ I Introduction
Recent advancements in the field of quantum machine learning (QML) Biamonte et al. (2017); Carleo et al. (2019); Dunjko and Briegel (2018) have significantly enhanced our understanding of how quantum resources can be utilized to design various new paradigms Huang et al. (2021a); Schuld (2021); Jerbi et al. (2023); Schuld and Killoran (2019). These developments have shed light on the trainability McClean et al. (2018); Cerezo et al. (2021); Holmes et al. (2022); Fontana et al. (2024); Ragone et al. (2023); Thanasilp et al. (2024); Pesah et al. (2021) and generalization capabilities of quantum models Caro et al. (2021, 2022); Peters and Schuld (2023); Banchi et al. (2021); Gil-Fuster et al. (2024) and have even demonstrated theoretical advantages over classical methods for certain tailored problems Gyurik and Dunjko (2023); Sweke et al. (2021); Jerbi et al. (2021); Liu et al. (2021); Pirnay et al. (2023). Despite these breakthroughs and the successful implementation of quantum models on quantum processors Havlíček et al. (2019); Johri et al. (2021); Ristè et al. (2017); Steinbrecher et al. (2019); Peters et al. (2021), the practicality of QML remains uncertain Jadhav et al. (2023).
Image classification is one of the most usfeul tasks due to its significant industrial applications. Various approaches have been explored, including quantum convolutional neural networks (QCNNs)Matic et al. (2022); Henderson et al. (2020a); Lü et al. (2021); Chen et al. (2023) and models that leverage symmetries Chang et al. (2023); West et al. (2023); Sebastian et al. (2024); West et al. (2024). There is increasing recognition that combining classical and quantum approaches can be more beneficial than trying to replace one by the other. The pursuit of better machine learning techniques with new methodologies involves a deeper consideration of various factors such as resource efficiency, data requirements, and the practical applications of these technologies. As addressed in this work, this involves pre-processing images classically before applying a quantum data processing model Senokosov et al. (2024); Dang et al. (2018). The aim is to demonstrate the development and implementation of models using components executable on quantum processors, with the goal of scaling them up to create models that cannot be simulated on classical processors.
@@
Not many studies have tackled this complex classification task using quantum resources. The prevalent approach involves using parameterized quantum circuits following dimensionality reduction techniques, as seen in Refs. Zollner et al. (2024); Sebastianelli et al. (2022); Benítez-Buenache and Portell-Montserrat (2024). Additionally, there have been approaches utilizing support vector machines (SVMs) based on quantum kernels Mountrakis et al. (2011); Delilbasic et al. (2021a). A common theme across these studies is the dependence on the choice of feature maps and/or ansatz, as well as the classical preprocessing methods used for dimensionality reduction or feature extraction prior to encoding. Building upon these foundations, the goal of this work is to address a practical use case with the potential to make a significant impact on the EO industry.
In this Article, we focus on the identification of solar photovoltaic (PV) panels, using a challenging real-world dataset Bradbury et al. (2016). This use case is motivated by the desire to contribute to a high-priority social and environmental application. Effectively addressing this problem has the potential to positively impact the logistics related to renewable energy industry, which is facing a significant increase in demand for renewable energy sources and the challenge of managing the escalating deployment of solar infrastructures worldwide. For this task, we present a comprehensive classical preprocessing pipeline for feature extraction based on a cross-task transfer learning implementation, which proves particularly effective for this purpose. To accommodate the data within the constraints of small-scale circuits in the Noisy Intermediate-Scale Quantum (NISQ) era Preskill (2018), dimensionality reduction techniques are employed. We construct two distinct quantum models based on methodologies outlined in Ref. Rodriguez-Grasa et al. (2024), both leveraging neural quantum kernels (NQKs), which are embedding quantum kernels (EQKs) obtained by training a quantum neural network (QNN).
The first model utilizes a $1$-to-$n$ NQK approach, where training a single-qubit QNN forms an $n$-qubit EQK. For this classification task, qubit numbers of $2$ and $3$ suffice. The second model employs an $n$-to-$n$ NQK strategy, using a trained $n$-qubit QNN directly as the embedding for constructing a quantum kernel. The training of the $n$-qubit QNN follows an iterative approach proposed in Ref. Rodriguez-Grasa et al. (2024), starting with a single qubit and scaling up to enhance classifier performance. Both approaches demonstrate robustness, achieving classification accuracies approaching 90%.
This paper is organized as follows: In Section II, we discuss related work, reviewing multiple references on quantum computing for EO, briefly analyzing the main trends in EO image classification, hybrid-QNNs, and Quantum Enhanced SVMs. Section III describes and summarizes the features and characteristics of the selected dataset, as well as the strategies adopted for pre-processing it. In Section IV, we introduce NQKs, a quantum classification method proposed by Rodriguez-Grasa et al. (2024). Section V presents and discusses the numerical results and their analysis. Finally, Section VI wraps up the manuscript by summarizing our findings, offering a critical assessment of the results, and suggesting directions for future research.
II Related work on satellite image classification
This section provides a brief overview of key contributions in remote sensing image classification using QML. We walk through few pertinent aspects, including feature extraction and dimensionality reduction, hybrid QNNs and Quantum SVMs (qSVMs).
II.1 Feature extraction and dimensionality reduction
The issue of dimensionality reduction and feature extraction is a common challenge in the context of machine learning. When dealing with image classification problems in remote sensing applications using QML, this issue becomes especially relevant. This is due to the high resolution of these images in question, which needs effective preliminary information processing and adaptation to fit into the limited number of qubits available in NISQ computers.
In Otgonbaatar and Datcu (2022), the problem of dimensionality reduction is addressed through the usage of a simple convolutional neural network (CNN) and an autoencoder in the first classic part of the solution pipeline working with the widely used RGB reduction of the EuroSAT image dataset (64x64 images crops from Sentinel 2 satellite images) Helber et al. (2019). With this approach the effective reduction of the dimensionality of the data is achieved, with the information ultimately being embedded in a 17-qubit parameterized quantum circuit (PQC). It is shown how the proposed PQC performs better than a fully connected dense layer and confirmed that the autoencoder is a better choice than the naive downsampling techniques.
On the same path, Zollner et al. (2024) compares different dimensionality reduction approaches for being adopted as preliminary classical preprocessing before using quantum circuits with different encoding and variational circuits to classify satellite images. Amongst dimensionality reduction techniques they consider downscaling, Principal Component Analysis (PCA), factor analysis, Restricted Boltzmann Machines (RBM), and both fully connected and convolutional autoencoders (CAE). They interestingly show how even smaller quantum systems have good potential for real applications, if applied with appropriate feature extraction techniques. As a matter of interest for our problem, they found the CAE based architectures being best suited as dimensionality reduction technique for extracting features from larger scale images.
In the field of dimensionality reduction techniques, the authors of Henderson et al. (2020b) explored the use of quantum convolutional filters learnt through direct training of a variational circuit applied to satellite images. They implemented an enhanced version of the Quanvolutional Neural Network for classifying the DeepSat SAT-4 dataset Basu et al. (2015), which was tested on real quantum hardware.
II.2 Hybrid Quantum Neural Networks
As illustrated in the above-described works, the concept of merging the flexibility of classical neural networks, and particularly CNNs, as an embedding mechanism with the expressivity power of variational quantum circuits is a recurring theme in the field of Quantum Computer for Earth Observation.
In Sebastianelli et al. (2022) and Zaidenberg et al. (2021), the authors demonstrate the application of hybrid QNNs to a multiclass land coverage classification problem. In their studies, they adopted the RGB EuroSAT dataset while introducing a PQC between the final fully connected layers of a classical CNN for classification. They highlighted the significance of the convolutional embedding step in encoding images prior to their use on quantum elements. They also explored the varying degrees of expressivity exhibited by alternative ansatz, including Bellman circuits (entanglement) and real amplitude circuits. The best model was the one based on the Real Amplitude circuit, with slightly inferior performance compared to the state of the art, but with a very small network (number of parameters) when compared with ResNet-based solutions. The approach was also validated with a sliding window approach on the Onera Satellite Change Detection Dataset (OSCD) Daudt et al. (2018).
The path for applying the concept of transfer learning to hybrid neural networks (quantum transfer learning) is explored in Otgonbaatar et al. (2023). The study focused on freezing parameters of the classical convolutional layers and only training the quantum layers of the architecture. They compared, against a classification task, several quantum circuits solutions and found that multi-qubit strongly entangled quantum layers perform better than real amplitude circuits in most experiments. Digging more into hybrid QNN for classification, Sebastianelli et al. (2023) opened the way to discussions about the complex topic of tuning of hyperparameters in hybrid QNNs. Using EuroSAT as a benchmark dataset, the authors proposed a practical strategy for identifying an efficient pair of values for the number of quantum layers (affecting the depth of the circuit) and the number of qubits (affecting the width of the circuit) in the PQC. This strategy uses contour plots representing the interpolation of several target metrics obtained through a number of initial experiments. Their work highlights the complexity and non-triviality of architecting and tuning a QNN for a specific task.
II.3 Quantum Support Vector Machines
In Delilbasic et al. (2021b), two distinct qSVM implementations were proposed, each leveraging a different main paradigm in quantum computing: gate-based circuits and quantum annealing. The authors demonstrated that these implementations for multispectral EO images are capable of producing comparable results with classic SVM in both simulation and real hardware. However, they also showed that they have clear limitations in terms of the amount of data that they can process on real hardware.
In the frame of qSVM it is interesting to mention the recent work from Ref. Centre. (2023) which verified that, while both quantum and classical runtimes scale logarithmically with dimensionality, quantum simulation times are larger than real quantum hardware ones: this finding observation leaves room for possible advantage with intermediate scale quantum computers.
@@ III.1 Data preparation
Figure 1: (a) Positive and (b) negative image samples with their binary segmentation masks
@@
    Image partitioning. Dealing with high-resolution images can be challenging in terms of memory consumption and computational time, even on classical hardware. In the NISQ era of quantum computing, this problem is further amplified due to the limited availability of sufficient qubits to encode data in quantum hardware and the great computational complexity required to implement large circuits in quantum simulators. To address this issue, each 5000-by-5000 pixel image is sliced into 400 smaller images of 250-by-250 pixels. This step allows for the creation of a dataset that includes both positive (panels present) and negative (panels absent) image samples. All of the original high-resolution images contain panels in them, but by cropping small sections of the images, it is possible to obtain samples that are empty of panels.
@@
    Sampling and rebalance classes. After slicing the original 463 images, there were 185,200 resulting sliced images, of which only 10,863 were positive instances, accounting for 5.768% of the total. Defining a “positive instance” for an image is not a straightforward task as it cannot be directly derived from the dataset. However, by using the image mask, it is possible to create a local criterion for assigning a class label. In this work, the class label for an image can be defined as
@@
    where $\gamma=\frac{1}{N}\sum_{i,j}M_{ij}$ is the positive pixel proportion for each image over the binary pixels of the mask $M_{ij}\in\{0,1\}$; $\varepsilon$ is a threshold to filter out “edge cases” and $N$ is the total number of pixels for each image. Note that if a given image does not comply with either of these conditions (i.e. if its mask contains a proportion of white pixels below the $\varepsilon$ threshold, but greater than zero), it will not be included in the dataset.
    Additionally, to facilitate quantum experimentation and simplify the problem, the dataset’s class imbalance was removed ad-hoc by undersampling the majority class.
@@
    Prune the positive cases. As defined in (1), it is necessary to establish a threshold to filter out non-representative cases. For instance, an image that contains only a few pixels of solar arrays can be indistinguishable from a negative image, even to the human eye. It is unrealistic to expect a classification model to accurately discern such “edge cases” where the object is not fully recognizable. Therefore, for simplicity, such cases are discarded from the dataset to avoid training the model with structures that could be confused with noisy background patterns.
    The value of the threshold $\varepsilon$ has been set based on the percentiles of the proportion of white pixels in the masks. For the entire dataset without filtering we define a percentile threshold $\varepsilon=T_{p}$ to filter out images with a low proportion of white pixels. The percentile rank $p$ was set heuristically to 15%, which corresponds to a presence of panels covering 0.2% of the image surface.
@@
Figure 2: Pipeline starting from the image up to obtaining the labels detailing the stages of classical preprocessing. Starting from $256\times 256\times 3$ images, the dimensionality is reduced to vectors of $p$ features ranging from -1 to 1.
Figure 3: Diagram illustrating the entire architecture from the input image to the NQK construction. The upper branch depicts the $1$-to-$n$ NQK, while the lower branch shows the $n$-to-$n$ NQK, built using an $n$-qubit QNN. This QNN will be trained iteratively, ensuring scalability.
III.2 Image pre-processing
In recent years, widespread methods for extracting features from high-dimensional (i.e. high-resolution) images have relied on the use of pre-trained classification CNNs (e.g., ResNet, EfficientNet, etc.) as well as the use of CAEs Ryu et al. (2020) . Based on some preliminary tests, the idea of solving the PV identification problem classically using pre-trained CNNs was discarded due to its low performance observed when benchmarking the extracted embeddings. This led to the conclusion that general purpose multiclass CNNs trained on ImageNet may not be suitable for detecting small objects in high resolution aerial or satellite images, which often contain various geometric structures and geospatial patterns. On the other hand, CAE architectures capable of learning to reconstruct the input image are often considered for feature extraction and dimensionality reduction of images before performing other downstream tasks. In both cases, the so-called bottleneck features (also known as the latent vectors) represent a global summary of the input image.
Given the specificity of the small structure identification problem in this work, we decide to approach the classification problem in a different way: from a “general to specific approach” to a “specific to general approach”. While identifying solar panels in an image is generally considered a global classification task, a more specific task is to identify the exact silhouette of the object in the image. Semantic segmentation techniques are commonly used to achieve this, using CAE architectures such as the well-known U-Net Ronneberger et al. (2015). Thus, it was determined that training a U-Net architecture from scratch using the images and their segmentation masks would yield optimal features for small object identification. This approach was previously proposed in Ref. Huang et al. (2021b) for classifying small lesions in medical images using features extracted from the bottleneck of semantic segmentation models.

### v2

@@ Abstract
Achieving practical applications of quantum machine learning for real-world scenarios remains challenging despite significant theoretical progress. This paper proposes a novel approach for classifying satellite images, a task of particular relevance to the earth observation (EO) industry, using quantum machine learning techniques. Specifically, we focus on classifying images that contain solar panels, addressing a complex real-world classification problem. Our approach begins with classical pre-processing to reduce the dimensionality of the satellite image dataset. We then apply neural quantum kernels (NQKs)—quantum kernels derived from trained quantum neural networks (QNNs)—for classification. We evaluate several strategies within this framework, demonstrating results that are competitive with the best classical methods. Key findings include the robustness of or results and their scalability, with successful performance achieved up to 8 qubits.
@@ I Introduction
Recent advancements in the field of quantum machine learning (QML) Biamonte et al. (2017); Carleo et al. (2019); Dunjko and Briegel (2018) have significantly enhanced our understanding of how quantum resources can be utilized to design various new paradigms Huang et al. (2021a); Schuld (2021); Jerbi et al. (2023); Schuld and Killoran (2019). These developments have shed light on the trainability McClean et al. (2018a); Cerezo et al. (2021a); Holmes et al. (2022a); Fontana et al. (2024); Ragone et al. (2023a); Thanasilp et al. (2024); Pesah et al. (2021) and generalization capabilities of quantum models Caro et al. (2021, 2022); Peters and Schuld (2023); Banchi et al. (2021); Gil-Fuster et al. (2024) and have even demonstrated theoretical advantages over classical methods for certain tailored problems Gyurik and Dunjko (2023); Sweke et al. (2021); Jerbi et al. (2021); Liu et al. (2021); Pirnay et al. (2023). Despite these breakthroughs and the successful implementation of quantum models on quantum processors Havlíček et al. (2019); Johri et al. (2021); Ristè et al. (2017); Steinbrecher et al. (2019); Peters et al. (2021), the practicality of QML remains uncertain Jadhav et al. (2023).
Image classification is one of the most usfeul tasks due to its significant industrial applications. Various approaches have been explored, including quantum convolutional neural networks (QCNNs)Matic et al. (2022); Henderson et al. (2020a); Lü et al. (2021); Chen et al. (2023); Simen et al. (2024); Altares-López et al. (2024a) and models that leverage symmetries Chang et al. (2023); West et al. (2023); Sebastian et al. (2024); West et al. (2024). There is increasing recognition that combining classical and quantum approaches can be more beneficial than trying to replace one by the other. The pursuit of better machine learning techniques with new methodologies involves a deeper consideration of various factors such as resource efficiency, data requirements, and the practical applications of these technologies. As addressed in this work, this involves pre-processing images classically before applying a quantum data processing model Senokosov et al. (2024); Dang et al. (2018). The aim is to demonstrate the development and implementation of models using components executable on quantum processors, with the goal of scaling them up to create models that cannot be simulated on classical processors.
@@
Few studies have addressed complex classification tasks using quantum resources; this work aims to bridge that gap by combining classical and quantum resources to demonstrate that quantum methods can achieve effective results in real-world scenarios. Our focus is on the identification of solar photovoltaic (PV) panels, utilizing a challenging dataset from a real-world scenario Bradbury et al. (2016) to address a high-priority social and environmental application. In order to achieve this, we developed an effective classical pipeline tailored to preprocess images from this complex satellite imagery dataset, reducing them to a limited number of features suitable for encoding in a quantum architecture. Building on previous studies that underscore the dependency of results on ansatz selection in approaches using parameterized quantum circuits Zollner et al. (2024); Sebastianelli et al. (2022); Benítez-Buenache and Portell-Montserrat (2024); Otgonbaatar and Datcu (2022) or on the choice of embedding in support vector machines based on quantum kernels Mountrakis et al. (2011); Delilbasic et al. (2021a, b); Centre. (2023); Altares-López et al. (2024b), we chose to employ neural quantum kernels (NQKs) Rodriguez-Grasa et al. (2024). This model uses a quantum neural network to generate a embedding quantum kernel (EQK) tailored to the classification task.
The decision to utilize NQKs addresses several challenges in the current quantum machine learning landscape. Firstly, it introduces a scalable model that mitigates trainability issues known in parametrized quantum circuits and quantum neural networks McClean et al. (2018b); Cerezo et al. (2021b); Holmes et al. (2022b); Anschuetz and Kiani (2022); Ragone et al. (2023b). As we show in this work, we achieve consistently improved results with the addition of qubits. Additionally, since training quantum neural networks can be problematic, constructing a kernel from the network has proven to be resilient to these training difficulties. This resilience allows us to demonstrate that constructing an effective kernel model does not require the ability to finely tune a quantum neural network. Finally, while the impact of entanglement on enhancing quantum machine learning models remained uncertain Bowles et al. (2024), our findings demonstrate that the performance improvements observed when adding qubits to neural quantum kernels are solely due to the presence of entanglement. This clarifies the crucial role of entanglement in these models.
Within the framework of NQKs, we consider two model constructions. The first, the $1$-to-$n$ model, is built by training a single-qubit QNN and then using the trained parameters to construct an $n$-qubit EQK. With this model, we observe that using just 2–3 qubits yields results comparable to classical counterparts, and we show its robustness with respect to the training of the single-qubit QNN. The second model, the $n$-to-$n$ model, involves training an $n$-qubit QNN by iteratively adding qubits to ensure scalability, then using the trained architecture to construct the corresponding $n$-qubit EQK. In this case, the trained QNN functions directly as the quantum embedding. We show that results improve progressively as qubits are added. With both models, we achieve accuracies approaching 90%.
This paper is organized as follows: Section III outlines the dataset generation process, explaining how $p$ features are extracted from the original images to serve as inputs for the quantum model. In Section IV, we introduce Neural Quantum Kernels (NQKs), a quantum classification approach proposed in Ref. Rodriguez-Grasa et al. (2024). Section V presents and analyzes the numerical results. Section VI explores the implications and limitations of our work, proposing directions for future research. Finally, Section VII summarizes our findings, providing a concise conclusion to the manuscript.
II Related Work
QML techniques have shown potential in advancing remote sensing image classification by addressing challenges like high-dimensional data and leveraging quantum computational power. This section provides an integrated overview of key contributions in feature extraction and dimensionality reduction, hybrid QNNs, and support vector machines based on quantum kernels.
Feature extraction and dimensionality reduction play a critical role in adapting high-resolution satellite images to the limited qubit capacity of NISQ computers. Studies such as Otgonbaatar and Datcu (2022) and Zollner et al. (2024) have demonstrated the effectiveness of classical preprocessing methods like convolutional autoencoders (CAE) and principal component analysis (PCA) for reducing dimensionality before quantum processing. Notably, CAE-based techniques were found to outperform naive downsampling methods, embedding relevant features into parameterized quantum circuits. Furthermore, Henderson et al. (2020b) introduced quantum convolutional filters trained through variational circuits, demonstrating their efficacy in classifying datasets like DeepSat SAT-4 using real quantum hardware.
Hybrid QNNs, combining classical neural networks with quantum circuits, have emerged as a promising approach for Earth observation. Works such as Sebastianelli et al. (2022) and Zaidenberg et al. (2021) utilized PQCs within classical convolutional pipelines to encode and classify RGB EuroSAT and Onera Satellite Change Detection datasets. These studies highlighted the importance of convolutional embeddings and explored various ansatz, with real amplitude circuits offering compact architectures but slightly lower performance compared to ResNet-based solutions. Additionally, Otgonbaatar et al. (2023) explored quantum transfer learning by freezing classical layers and training quantum layers, finding that strongly entangled multi-qubit circuits performed best. Hyperparameter tuning in hybrid QNNs, a complex yet critical aspect, was systematically addressed in Sebastianelli et al. (2023), providing strategies for optimizing circuit depth and qubit count using EuroSAT as a benchmark.
Various studies have explored the use of quantum kernels for solving real-world tasks, though these kernels are typically problem-agnostic Huang et al. (2021b); Peters et al. (2021); Bartkiewicz et al. (2020); Kusumoto et al. (2021); Wu et al. (2023); Kyriienko and Magnusson (2022). The development of neural quantum kernels, however, opens the door to creating quantum kernels specifically tailored to individual problems. In the context of satellite image classification, both gate-based circuits and quantum annealing have been applied, as demonstrated in Delilbasic et al. (2021b), where these methods achieved performance comparable to classical SVMs on multispectral data. Despite this, real hardware limitations imposed constraints on data scalability. Recent advancements, such as those in Centre. (2023), revealed that while both classical and quantum runtimes scale logarithmically with dimensionality, real quantum hardware can offer runtime advantages over simulations, suggesting potential for intermediate-scale quantum systems.
Collectively, these works underline the synergistic potential of combining classical preprocessing with quantum methodologies, while also highlighting the challenges and opportunities in developing scalable and efficient QML frameworks for remote sensing applications.
@@ III The Dataset
The following subsections, III.1 and III.2, report details about the dataset preparation and feature engineering steps, required for the following implementation of classical and quantum classification models. The whole workflow is shown in Figure 2.
@@ III.1 Data preparation
Figure 1: (a) Positive image sample with solar panels ($\hat{y}=+1$) and (b) negative image sample without solar panels ($\hat{y}=-1$), each accompanied by their respective binary segmentation masks. The negative sample illustrates a case where the proportion of white pixels is $\gamma=0$.
@@
    Image partitioning. Working with high-resolution images is demanding in terms of memory usage and computational time. This challenge, already significant for classical hardware, becomes even more pronounced on quantum devices due to limited quantum resources, such as the number of available qubits and allowable circuit depth. To address these constraints, each 5000-by-5000-pixel image is divided into 400 smaller images, each measuring 250-by-250 pixels. While all original high-resolution images contain panels, cropping small sections of these images produces samples without panels, enabling the creation of a dataset with both positive samples (images containing panels) and negative samples (images without panels).
@@
    Sampling and rebalance classes. After slicing the original 463 images, we obtained 185,200 smaller images, of which only 10,863 were positive instances, representing 5.768% of the total. Defining a “positive instance” for each image is not straightforward, as this information is not directly available in the dataset. However, using the image mask, we can establish a local criterion to assign class labels. In this work, an image’s class label is defined as follows
@@
    where $\gamma=\frac{1}{N}\sum_{i,j}M_{ij}$ represents the proportion of positive (white) pixels in each image mask $M_{ij}\in\{0,1\}$, $\varepsilon$ is a threshold set to exclude “edge cases” and $N$ is the total number of pixels in each image. Note that if an image’s mask contains a proportion of white pixels between zero and the threshold $\varepsilon$, it is excluded from the dataset.
    Additionally the dataset’s class imbalance was removed ad-hoc by undersampling the majority class.
@@
    Prune the positive cases. As defined in Eq. 1, it is necessary to establish a threshold to filter out non-representative cases. For instance, an image containing only a few pixels of solar arrays may be visually indistinguishable from a negative image, even to the human eye. Expecting a classification model to accurately identify such ”edge cases,” where the object is barely recognizable, would be unrealistic. Therefore, to avoid training the model on structures that might be confused with background noise, these ambiguous cases are excluded from the dataset.
    The threshold value $\varepsilon$ for filtering was determined based on the distribution of solar array pixels across all cropped images. Specifically, the 15th percentile of the pixel count distribution was selected as the minimum value to retain positive samples. This threshold corresponds to a maximum panel coverage of 0.2% of the entire cropped image area and balances the goal of retaining sufficient samples while ensuring a high-quality dataset.
@@
Figure 2: Flowchart of the image processing workflow. The input dataset for the classification model is prepared, and features are extracted to meet the requirements for encoding the information into a quantum model.
Figure 3: Diagram illustrating the complete architecture, from the input image to the construction of the NQK. Starting with large, complex images, a dataset $X$ of $M$ data points is generated, consisting of 2- or 3-dimensional points, through classical preprocessing. The upper branch represents the $1$-to-$n$ NQK construction, while the lower branch illustrates the $n$-to-$n$ NQK, which is built using an $n$-qubit QNN. This QNN is trained iteratively to ensure scalability. $\vec{\theta}$ and $\vec{\varphi}$ denote the trainable parameters.
III.2 Feature engineering
Since the input to the problem consists of geospatial images, and the purpose of this work is to study QML algorithms designed for tabular data, it becomes necessary to extract tabular features from the images. In recent years, widespread methods for extracting features from high-dimensional (i.e. high-resolution) images have relied on the use of pre-trained classification CNNs (e.g., ResNet, EfficientNet, etc.) as well as the use of CAEs Ryu et al. (2020) . Preliminary tests indicated that solving the PV identification problem classically using pre-trained CNNs was not viable, as benchmarking the extracted embeddings revealed low performance. This observation led to the conclusion that general-purpose multiclass CNNs trained on ImageNet may be unsuitable for detecting small objects in high-resolution aerial or satellite images, which typically feature diverse geometric structures and geospatial patterns. On the other hand, CAE architectures capable of learning to reconstruct the input image are often considered for feature extraction and dimensionality reduction of images before performing other downstream tasks. In both cases, the so-called bottleneck features (also known as the latent vectors) represent a global summary of the input image.
Given the specificity of the small structure identification problem in this work, we decide to approach the classification problem in a different way: from a “general to specific approach” to a “specific to general approach”. While identifying solar panels in an image is generally considered a global classification task, a more specific task is to identify the exact silhouette of the object in the image. Semantic segmentation techniques are commonly used to achieve this, using CAE architectures such as the well-known U-Net Ronneberger et al. (2015). Thus, it was determined that training a U-Net architecture from scratch using the images and their segmentation masks would yield optimal features for small object identification. This approach was previously proposed in Ref. Huang et al. (2021c) for classifying small lesions in medical images using features extracted from the bottleneck of semantic segmentation models.